OpenStudy (anonymous):

If r= 25 & y= 15, then x= ?

OpenStudy (lgbasallote):

\[x^2 + y^2 = r^2\] \[x^2 = r^2 - y^2\] \[x = \sqrt{r^2 - y^2}\] \[x = \sqrt{25^2 - 15^2}\]

OpenStudy (anonymous):

20?

jimthompson5910 (jim_thompson5910):

you are correct

OpenStudy (anonymous):

if x= 10 and y= 6 what is r? how would i set that up?>

jimthompson5910 (jim_thompson5910):

\[\Large x^2 + y^2 = r^2\] \[\Large (10)^2 + (6)^2 = r^2\] \[\Large 100 + 36 = r^2\] I'll let you finish

OpenStudy (anonymous):

i got 4 square root 17 ?

jimthompson5910 (jim_thompson5910):

\[\Large 100 + 36 = r^2\] \[\Large 136 = r^2\] \[\Large r^2 = 136\] \[\Large r = \sqrt{136}\] \[\Large r = \sqrt{4*34}\] \[\Large r = \sqrt{4}*\sqrt{34}\] \[\Large r = 2\sqrt{34}\]

jimthompson5910 (jim_thompson5910):

So you were close

jimthompson5910 (jim_thompson5910):

somehow you factored out an additional 2